OpenStudy (anonymous):

Multiply (4 - sqrrt of 3)(12 + 5(sqrrt of 3)). Then simplify.

OpenStudy (schrodinger):

\[(4 - \sqrt{3})(12 + 5\sqrt{3})\]From here, just use FOIL: \[4(12) + 4(5\sqrt{3}) - \sqrt{3}(12) - 5\sqrt{3}(\sqrt{3})\]

OpenStudy (schrodinger):

\[48 + 20\sqrt{3} - 12\sqrt{3} - 15\]

OpenStudy (schrodinger):

\[33 + 8\sqrt{3} \] ??? (I'm pretty sure that's it, but i'm not great with math so i'm not willing to say it ends there. That's what I can say, though.

OpenStudy (anonymous):

Thank you :D You really helped.